OpenStudy (anonymous):

According to Josh’s measurement, the distance from his house to the baseball practice field is 4 km. The actual distance is 3.68 km. What was the percent error in Josh’s measurement? Round to the nearest tenth of a percent, if needed. A. 0.3% B. 8% C. 8.7% D. 32% PLZ HELP I WILL GIVE OUT MEDAL!!!

OpenStudy (anonymous):

i think B

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

@pooja195 @PlanSlam @weirdawesomegirl10 @KendrickLamar2014 @mathslover

OpenStudy (anonymous):

plz help me

OpenStudy (anonymous):

yup i got B

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

thx @PlanSlam

OpenStudy (anonymous):

how did u get b

OpenStudy (anonymous):

i am just askin

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

@PlanSlam how did u get b i am just askin so i can know @sleepyjess

OpenStudy (anonymous):

Ok this problem works like this. I am going to use counting numbers, which makes it simpler I think. Lets say we measured 4 but the actual measurement was 2 . First fine the difference 4 - 2 = 2 Now what percent of 4 is 2. This is like saying what fraction of 4 is 2 2 / 4 = 1/2 = .5 = 50% The problem you have works the same way 4 - 3.68 = .32 Can you finish the rest?

OpenStudy (anonymous):

@OpenStudyBot

OpenStudy (anonymous):

u r the best thx i am 99.9998999989% sure that i got it know

OpenStudy (anonymous):

now

OpenStudy (openstudybot):

It is B.